The Soma Inn is trying to determine its break-even point. The inn has 75 rooms that are rented at $52 a night. Operating costs are as follows.

Salaries $4,900 per month
Utilities 2,500 per month
Depreciation 1,500 per month
Maintenance 850 per month
Maid service 7 per room
Other costs 32 per room

Determine the inn’s break-even point in (1) number of rented rooms per month and (2) dollars.

1. Break-even point in rooms
2. Break-even point

$

per month
Answer #1

Total variable costs=Maid service+Other costs

=7+32=$39 per room

Total fixed cost=Salaries+Utilities+Depreciation+Maintenance

=4900+2500+1500+850=$9750

Contribution margin=Sales-Variable cost  

=(52-39)=$13 per room

Breakeven=Fixed expenses/Contribution margin

=(9750/13)=750 rooms

=(750*52)=$39000